Sarah Bates, calendar year taxpayer, started a new business on October 8th. A number of start-up expenditures (wages, utilities, rent, etc) were paid before the business offically opened. What is the deduction for these expenditures for the first year?. (Do not round intermediate calculations. Round your final answers to the nearest whole dollar amount.)

  1. The start-up expenditures totaled $4,910.
  2. The start-up expenditures totaled $33,080.
  3. The start-up expenditures totaled $52,800.
  4. The start-up expenditures totaled $94,140.

a.Deduction:

b.Deduction:

c.Deduction:

d.Deduction :
Answer #1

For start up costs $5,000 is written off in the first year. Reduce $ for $ for total costs greater than $50,000 and amortize the balance over 180 months including the first year from start up month.

a. Deduction = $4,910

b. Deduction = $5,000 + ($33,080 - $5,000) / 180 * 3 = $5,468

c. Deduction = $2,200 + ($52,800 - $2,200) / 180 * 3 = $3,043

d. Deduction = $94,140 / 180 * 3 = $1,569